OpenStudy (anonymous):

A population of flies grows according to the function p(x) = 2(4)x, where x is measured in weeks. A local spider has set up shop and consumes flies according to the function s(x) = 2x + 5. What is the population of flies after two weeks with the introduced spider? 15 flies 23 flies 32 flies 36 flies

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Princezz

OpenStudy (anonymous):

What is the first function?

OpenStudy (anonymous):

p(x) = 2(4)

OpenStudy (anonymous):

what that means? Is it p(x) = 2*4 = 8 ?